Problem:
The standard deviation of the breaking strengths of certain cables produced by a company is given as 240 lbs. After a change was introduced in the process of manufacture of these cables, the breaking strengths of a sample of 8 cables showed a standard deviation of 300 lbs. investigate the significance of the apparent increase in variability, using a significance level of (a) 0.05 and (b)0.01.
1
Expert's answer
2020-05-15T17:43:53-0400

σ0​=240, N=8, s=300

We assume that the breaking strengths has normal distribution

a)α=0.05

H0​:σ2=σ02​=2402,H1​:σ2​≠ σ02​=2402

x2=((N-1)s2)/σ02, s2 is a random value of corrected variance

This random variable has x2distribution with k=N−1=7 degree of freedom.

Observed value:x2≈10.9375.Critical values: x2crleft=x2cr(1-α/2;k)=1.6899

x2crright=x2cr(α/2;k)=16.013

Critical region(-∞;1.6899)"\bigcup" (16.013;)

Our observed value does not fall into the critical region.So we accept H0 - Variability did not change after a change in the process of manufacture.

b)α=0.01

Critical values: x2crleft=x2cr(1-α/2;k)=0.9893

x2crright=x2cr(α/2;k)=20.278

Critical region(-∞;0.9893)(20.278;)

Our critical value does not fall into the critical region.So we accept H0 Variability did not change after a change in the process of manufacture.